Emily has $126 in her checking account. She writes a check for $113, makes a deposite of $50, and then writes another check for $100. What is the balance of her account, written as an integer?
Answer by nyc_function(2741)

Emily's total is $126.
She writes a check for 113 and makes a deposite of 50.
So, 126 - 113 = 13 + a deposite of 50 = 63.
Then she decides to write another check for 100.
This was a big mistake because she only has $63 in the bank.
This means that 63 - 100 = -37.
Her balance is -37 dollars in the bank which is not a good account at all.
Emily is in trouble.